What is the present value of $4,000 per month for one year with an annual interest rate of 7.2% (APR with monthly compounding)?
Answer #1

APR = 7.2%

monthly rate = 7.2%/12 =0.6% = 0.006 = r

no of months = 12 = n

PMT = 4000

PV = PMT*(1/r)*(1-1/(1+r)^n))

PV = 4000*(1/0.006)*(1-1/(1+0.006)^12))

PV = 4000*166.6667*0.06926888 = 46179.26

Answer : 46179.26 [Thumbs up please]